Equilibrium is the condition in the market system when the quantity demanded is equal to the quantity supplied. Given the definition of equilibrium, answer the following questions:
4.2.1. Calculate the equilibrium quantity and price if the quantity supplied can be represented by the equation Qs = 18 000 + 0.2P and the quantity demanded can be represented by the equation Qd = 2 400 – 0.1P. [Tip: Solve the value of P first and then substitute this P-value in the calculation of the Q-value.]
1
Expert's answer
2021-03-30T19:59:39-0400

Qs=qd

18000+0.2p=24000-0.1p

24000-18000=0.2p+0.1p

6000=0.3p

P=6000/0.3

P=20000

Qd=24000-0.1p

Qd=24000-0.1(20000)

Qd=22000